OpenStudy (anonymous):

the supp of an angle is 60 less than the twice the supp. of the comp angle. find the measure of the comp supp = supplementary comp = complementary

OpenStudy (anonymous):

Could you fill in 90 for complementary and 180 for supplementary, and you would then have to divide and subtract(maybe) to get your answer.... I think... try it and see what you get!

OpenStudy (anonymous):

i know the 1st part is 180-x=60_2x then i dont no the last part

jimthompson5910 (jim_thompson5910):

let x = unknown angle the supplement of this angle is 180 - x the complement of this angle is 90-x ------------------------------------------------------- "the supp of an angle is 60 less than the twice the supp. of the comp angle." translates to 180 - x = 2*(90-x) - 60 solve for x and tell me what you get

OpenStudy (anonymous):

wait wheres the supp on the right of the equal sign

jimthompson5910 (jim_thompson5910):

oh wait, i see what you mean

jimthompson5910 (jim_thompson5910):

let x = unknown angle the supplement of this angle is 180 - x the complement of this angle is 90-x ----------------------------------- the supplement of the complement is 180 - (90 - x) 180 - 90 + x 90 + x ---------------------------------- so "the supp of an angle is 60 less than the twice the supp. of the comp angle." translates to 180 - x = 2*(90 + x) - 60

OpenStudy (anonymous):

ok thanks

jimthompson5910 (jim_thompson5910):

np

jimthompson5910 (jim_thompson5910):

what did you get up getting for the answer